OpenStudy (anonymous):

Abigal Orders a large 2 topping pizza for $14 and a medium 1 topping pizza for $10. If all costs of a medium pizza are 80% of a similar large pizza, find the price of a large plain cheese pizza and the cost of each larger izza topping

OpenStudy (anonymous):

large 1 topping pizza = 10/8 * 10 = 12.5 price of each topping = large 2 topping - large 1 topping = 14 - 12.5 = 1.5 price of large 0 topping = 12.5 - 1.5 = 11